OpenStudy (anonymous):

whats 3 divided my radical 3

OpenStudy (anonymous):

Assuming you mean "by radical 3" and not "my radical 3", you should get just radical 3. Because you can't properly divide by a radical, you multiply the numerator and denominator by said radical. Thus, leaving you with\[\frac{ 3\sqrt{3} }{ 3 }\] and the 3's cancel, leaving you with just radical 3.

OpenStudy (anonymous):

it is always true that \[\frac{a}{\sqrt{a}}=\sqrt{a}\]